prettygirl15:

Pythagorean Theorem - a2 + b2 = c2 12 right-triangle-3.svg 20 20. Pythagorean Theorem - a2 + b2 = c2 Solve with pythagorean theorem where a= 5, and b= 10.

Joe348:

yo u got a pic?

THEKINGOFTHEMATRIX:

Did she add one? No.

prettygirl15:

i got it

THEKINGOFTHEMATRIX:

"Don't assume stuff buddy"

Joe348:

well can you screenshot it please and post

Joe348:

no?

THEKINGOFTHEMATRIX:

😂😂😂😂😂😂😂

Vocaloid:

a^2 + b^2 = c^2 the problem gives you a = 5 and b = 10, so plug these in and solve for c

prettygirl15:

@joe348 wrote:
yo u got a pic?
who